An engineer has designed a valve that will regulate water pressure on an automobile engine. The valve was tested on 220 engines and the mean pressure was 5.2 lbs/square inch. Assume the standard deviation is known to be 1. If the valve was designed to produce a mean pressure of 5.1 lbs/square inch, is there sufficient evidence at the 0.05 level that the valve performs above the specifications?

State the null and alternative hypotheses for the above scenario.

Please show me how you got the answer! Thank you!
Answer #1

Solution :

Given that,

Population mean = \mu = 5.1

Sample mean = \bar x = 5.2

Population standard deviation = \sigma = 1

Sample size = n = 220

Level of significance = \alpha = 0.05

This a right (One) tailed test.

The null and alternative hypothesis is,  

Ho: \mu \leq 5.1

Ha: \mu > 5.1

The test statistics,

Z =( \bar x - \mu )/ (\sigma/\sqrt{}n)

= ( 5.2 - 5.1 ) / ( 1 / \sqrt{} 220 )

= 1.48

P-value = P( Z > z )

= 1 - P(Z < 1.48 )

= 0.0694

The p-value is p = 0.0694, and since p = 0.0694 > 0.05, it is concluded that the null hypothesis is fail to rejected.

There is not sufficient evidence at the 0.05 level that the valve performs above the specifications.